OpenStudy (colcaps):

does a convex nonagon has 27 diagonals?

ganeshie8 (ganeshie8):

nonagon = how many sides ?

OpenStudy (colcaps):

9=nona?

OpenStudy (colcaps):

but what does convex mean?

ganeshie8 (ganeshie8):

convex means all interior angles are < 180

ganeshie8 (ganeshie8):

i think we dont have to use that info, convex/concave, it will have same # of diagonals

OpenStudy (colcaps):

so it has 27 diagonals?

ganeshie8 (ganeshie8):

did u use this formula : 9(9-3)/2

ganeshie8 (ganeshie8):

?

OpenStudy (colcaps):

damn geometry is hard sorry if your having a hard time with me

ganeshie8 (ganeshie8):

noo im enjoying doing these wid u, actually i like geometry and i did these very long time back, now its nice refresh :)

OpenStudy (colcaps):

so 54/2=27

ganeshie8 (ganeshie8):

thats right !

OpenStudy (colcaps):

i have another question the perimeter of a regular decagon is 200 cm. Each side measures 20 cm

ganeshie8 (ganeshie8):

regular decagon means, all 10 sides are congruent

OpenStudy (colcaps):

so 10/200? or should i use the formula?

ganeshie8 (ganeshie8):

since each side is 20cm, Perimeter = 10 * S = ?

OpenStudy (colcaps):

so its 200 10x20

ganeshie8 (ganeshie8):

10/200 doesnt give meaning ful info,

ganeshie8 (ganeshie8):

thats right :) since one side is 20cm 10 sides will be 10*20 = 200 so perimeter = 200

OpenStudy (colcaps):

but it can still give the answer?

ganeshie8 (ganeshie8):

this question is a True/False quesiton right ?

ganeshie8 (ganeshie8):

it doesnt ask us to calculate anything,

OpenStudy (colcaps):

yes

OpenStudy (colcaps):

so its true

OpenStudy (colcaps):

can i ask more questions?

ganeshie8 (ganeshie8):

sure ask :)

OpenStudy (colcaps):

ill post another question ill close this

ganeshie8 (ganeshie8):

k